OpenStudy (mathsolver):

Given that log6=log3+log2, express log5 in terms of log2 and log3

OpenStudy (anonymous):

don't log5 is breakable into 2 and 3, since 5 is a prime.

OpenStudy (anonymous):

think*

OpenStudy (mathsolver):

That's what I thought :( that was pulled straight from an exam paper though. Do you have any idea how to solve it?

OpenStudy (anonymous):

no idea

OpenStudy (raden):

log 5 = log (10/2) = log 10 - log 2 = 1 - log 2 = (log 3/log 3) - log 2